### v1

# Equi-RO: A 4D mmWave Radar Odometry via Equivariant Networks

Abstract
Autonomous vehicles and robots rely on accurate odometry estimation in GPS-denied environments. While LiDARs and cameras struggle under extreme weather, 4D mmWave radar emerges as a robust alternative with all-weather operability and velocity measurement. In this paper, we introduce Equi-RO, an equivariant network-based framework for 4D radar odometry. Our algorithm pre-processes Doppler velocity into invariant node and edge features in the graph, and employs separate networks for equivariant and invariant feature processing. A graph-based architecture enhances feature aggregation in sparse radar data, improving inter-frame correspondence. Experiments on the open-source dataset and self-collected dataset show Equi-RO outperforms state-of-the-art algorithms in accuracy and robustness. Overall, our method achieves 10.7% and 20.0% relative improvements in translation and rotation accuracy, respectively, compared to the best baseline on the open-source dataset.
I Introduction
Odometry is indispensable for autonomous vehicles and robots, providing accurate support for downstream perception, planning, and control. When Global Positioning System (GPS) signals are degraded, such as under bridges, in tunnels, or during severe rain or snow, it becomes crucial to obtain relative localization using sensors such as LiDARs and cameras.
However, the performance of LiDARs and cameras degrades significantly under extreme weather conditions, limiting their applicability for odometry estimation in such scenarios. Recently, 4D millimeter-wave (mmWave) radar—characterized by compact size, cost efficiency, all-weather adaptability, velocity-measuring capability, and long detection range—has been increasingly regarded as a promising solution for robust perception and localization in autonomous driving, particularly under adverse weather conditions [1].
Current 4D mmWave radar odometry algorithms are mostly adapted from traditional LiDAR odometry methods. Some exploit unique properties of 4D mmWave radar point clouds like Doppler velocity and Radar Cross Section (RCS) to perform ego-velocity estimation and point cloud registration [2, 3]. Regarding learning-based approaches, a few pioneering works have adapted Convolutional Neural Networks (CNNs) and Recurrent Neural Networks (RNNs) for stand-alone 4D mmWave radar [4] and radar-camera fusion [5] odometry.
Figure 1: Our Equi-RO method incorporates Doppler velocity compensation and graph construction to preserve noise-resilient and rotation-robust features. The derived equivariant and invariant features are fed into an equivariant Graph Neural Network (GNN)-based framework to obtain consistent and generalizable point cloud registration, yielding more accurate odometry results, especially in large rotations. Trajectory comparison on a partial segment of the loop3 split from the NTU4DRadLM dataset[6], with initialization aligned to the ground truth at the segment’s start, against [2].
In learning-based odometry, accurate feature extraction and matching are critical steps. However, features in adjacent frames may become highly inconsistent when the ego-vehicle moves rapidly or turns sharply. Equivariant neural networks have emerged as a promising approach to address this issue [7, 8]. They learn features that transform predictably under specific geometric transformations, such as 3D translations and rotations represented by rigid-body motion group SE(3), thereby enhancing feature matching performance. The geometric symmetry can improve the generalizability and interpretability of learning-based odometry systems.
The 4D mmWave radar provides additional physical attributes compared to LiDAR point clouds. For example, it measures the RCS of each point, which can be regarded as SE(3)-invariant since it reflects the strength of the returned signal and is determined mainly by the target’s size, shape, and material rather than the sensor’s pose. In addition, the Doppler effect enables estimation of the radial relative velocity of detected targets. However, Doppler velocity is neither equivariant nor invariant under SE(3) transformations, as it depends on both the sensor’s motion and the target’s motion, which limits the applicability of existing equivariant networks to 4D mmWave radar odometry. Moreover, 4D mmWave radar point clouds are inherently noisier, sparser, and more irregular than those captured by scanning LiDARs. Consequently, effectively aggregating point-wise features with those of their spatial neighbors remains a significant challenge for learning-based 4D mmWave radar odometry.
In this paper, we propose Equi-RO, a novel framework for 4D mmWave radar odometry built upon equivariant neural networks. Compared with existing 4D mmWave radar odometry algorithms, we design a pre-processing pipeline to derive invariant Doppler features and construct a graph representation to mitigate noise and sparsity. The equivariant and invariant features are then processed through separate networks to estimate the relative transformation. The invariant velocity features and equivariant network design enable our method to remain robust in challenging real-world driving scenarios, where existing methods often degrade. Leveraging both a public dataset [6] and a self-collected dataset, we demonstrate consistent improvements over state-of-the-art 4D mmWave radar odometry approaches.
In summary, our main contributions are as follows:
  1. 1.
    We develop a novel 4D mmWave radar odometry algorithm based on equivariant neural networks, capable of producing accurate and robust odometry estimates under large rotations and fast motions, as commonly encountered in sharp turns and challenging driving maneuvers.
  2. 2.
    We design a dedicated pre-processing pipeline for Doppler velocity that derives invariant features, thereby stabilizing motion-related features. And we design a framework that separately extracts equivariant and invariant features to preserve geometric consistency, making it applicable to 4D mmWave radar point clouds with physical attributes.
  3. 3.
    We conduct extensive experiments on both an open-source dataset and a self-collected dataset, along with ablation studies, to validate the effectiveness of the proposed method across diverse scenarios.
  4. 4.
    The source code will be released after receiving the final decision.
II Related Works
In this section, we briefly review existing algorithms for 4D mmWave radar odometry, followed by a discussion of equivariant neural networks.
II-A 4D mmWave Radar Odometry
Existing 4D mmWave radar odometry approaches can be broadly categorized into traditional and learning-based methods.
Traditional odometry methods, which do not involve neural networks, primarily address two challenges: (1) the effective utilization of Doppler velocity and RCS information, and (2) the handling of noise and sparsity in radar point clouds. For the first aspect, Doppler velocity is commonly employed for ego-velocity estimation [9] and dynamic object removal [3], while RCS is often used for point cloud filtering [10] and weighted point matching [11]. In the latter case, some works have proposed to use scan-to-submap registration [3] and adaptive probability-based registration [2, 12] to handle the noise and sparsity.
Learning-based approaches replace traditional point cloud registration and pose regression with deep neural networks. As an early pioneer, Lu et al. [4] employed CNNs to extract features from radar point clouds. RaFlow [13] formulates odometry estimation by scene flow prediction. 4DRO-Net [14] and 4DRVO-Net [5] integrate velocity-aware attention cost volumes into a coarse-to-fine optimization framework to iteratively estimate poses. CAO-RONet [15] incorporates local completion and context-aware association to match noisy points and suppress outliers. To the best of our knowledge, although GNNs can explicitly aggregate features from spatially neighboring points, they have not yet been applied to 4D mmWave radar odometry.
II-B Equivariant Neural Networks
Equivariant neural networks are designed to produce predictable changes in their outputs when specific transformations are applied to the inputs. This property enhances both the generalizability and interpretability of the models. Since Cohen and Welling first introduced the concept of equivariance into neural networks [7], various equivariant designs have been proposed for GNNs [16, 17]. For example, Satorras et al. [17] propose Equivariant Graph Neural Network (EGNN), which maintains equivariance to the Euclidean group E(n)—including translations, rotations, and reflections—by incorporating the relative squared distances between coordinates in the edge feature aggregation process.
Significant efforts have also been devoted to designing equivariant architectures for point cloud analysis. Tensor Field Network (TFN) [18] employs spherical harmonics to construct learnable weight kernels, preserving equivariance to SE(3) for point cloud data. Vector Neurons (VN) [19] and Lie Neurons (LN) [20] extend traditional Multi-Layer Perceptrons (MLPs) to operate on vector features, achieving equivariance to 3D rotation group SO(3) and semisimple Lie groups, respectively. EPN [21] and E2PN [8] further introduce SE(3)-equivariant convolutional architectures for 3D point clouds. Despite these advances, the application of equivariant neural networks to 4D mmWave radar point clouds remains largely unexplored.
III Methodology
This section provides a detailed description of our proposed Equi-RO algorithm. We first present an overview of the approach in Sec. III-A, followed by the pre-processing pipeline for 4D mmWave radar point clouds in Sec. III-B. Sec. III-C introduces the graph construction process and the definition of node and edge features, while Sec. III-D outlines the overall network architecture of Equi-RO.
III-A Overview
The overall pipeline of Equi-RO is illustrated in Fig. 2. Two consecutive frames of 4D mmWave radar point clouds are first pre-processed to estimate the ego-vehicle’s velocity and to generate invariant compensated node and edge velocities based on Doppler velocity. The processed point clouds are represented as graphs, with carefully defined node and edge features. An SE(3)-equivariant network based on LN [20] is employed to extract and aggregate node-level equivariant features, while both node- and edge-level invariant features are fed into an EGNN-based [17] module. The combined features are ultimately used to estimate the relative transformation $(R,t)$ between the two input frames, achieving robust registration even under large rotations.
Figure 2: Overview of the Equi-RO algorithm. Doppler velocity is used to derive invariant node and edge features, which are combined with node and edge features within a unified graph-based network framework to robustly estimate the relative transformation $(R,t)$ between two consecutive radar frames.
III-B Point Cloud Pre-processing
As mentioned above, 4D mmWave radar can directly measure the radial relative velocity of detected targets via the Doppler effect. In this section, we fully exploit the Doppler velocity to estimate the ego-vehicle’s velocity and to derive compensated node and edge velocities for subsequent use in the GNN-based equivariant framework.
III-B1 Ego-velocity Estimation
We apply an iterative reweighted least squares algorithm [22] to estimate the ego-velocity $\mathbf{v}_{ego}$. The Doppler velocity $v_{i}^{dop}$ of each point $i$ is affected by the coordinate $\bm{\rho}_{i}$ of this point in the sensor frame. In this work, Doppler velocity is defined as positive for approaching targets and negative for receding ones. For a point with absolute velocity $\mathbf{v}_{i}^{abs}$, its Doppler velocity $v_{i}^{dop}$ can be represented as:
$v_{i}^{dop}=\frac{\bm{\rho}_{i}^{T}}{\left\|\bm{\rho}_{i}\right\|}\cdot(\mathbf{v}_{ego}-\mathbf{v}_{i}^{abs}),$ (1)
where $\bm{\rho}_{i}^{T}$ is the transpose of $\bm{\rho}_{i}$. Assuming that most points in the scene are stationary, and assigning each point $i$ a weight $\omega_{i}$ indicating its likelihood of being stationary, we formulate the following weighted least squares problem for a point cloud with $N$ points:
$\underset{\mathbf{v}_{ego}}{\min}\sum_{i=1}^{N}\omega_{i}\left\|v_{i}^{dop}-\frac{\bm{\rho}_{i}^{T}}{\left\|\bm{\rho}_{i}\right\|}\cdot\mathbf{v}_{ego}\right\|.$ (2)
Here, the weights are iteratively updated as $\omega_{i}=1/(v_{r}+\epsilon)$, where $v_{r}=\left\|v_{i}^{dop}-\frac{\bm{\rho}_{i}^{T}}{\left\|\bm{\rho}_{i}\right\|}\cdot\mathbf{v}_{ego}\right\|$ and $\epsilon=1\times 10^{-5}$. Upon convergence, $\mathbf{v}_{ego}$ provides a robust estimate of the ego-vehicle’s velocity.
III-B2 Node Velocity Compensation
Once $\mathbf{v}_{ego}$ is estimated, the compensated node velocity $v_{i}^{dop^{\prime}}$ for each point $i$ can be computed as:
$v_{i}^{dop^{{}^{\prime}}}=v_{i}^{dop}-\frac{\bm{\rho}_{i}^{T}}{\left\|\bm{\rho}_{i}\right\|}\cdot\mathbf{v}_{ego}.$ (3)
As illustrated in Fig. 3, $v_{i}^{dop^{\prime}}$ corresponds to $\frac{\bm{\rho}_{i}^{T}}{\left\|\bm{\rho}_{i}\right\|}\cdot\mathbf{v}_{i}^{abs}$, i.e., the radial component of the absolute velocity $\mathbf{v}_{i}^{abs}$ of point $i$. Notably, $v_{i}^{dop^{\prime}}$ is invariant under SO(3) transformations of the radar, providing a stable feature representation for point cloud registration under large rotational variations.
Since $\mathbf{v}_{ego}$ has already been obtained to enhance relative translation estimation, incorporating the SO(3)-invariant $v_{i}^{dop^{\prime}}$ further strengthens the robustness of feature matching, thereby improving rotational alignment accuracy.
III-B3 Edge Velocity Compensation
A graph is constructed from the radar point cloud, so it is necessary to define an edge feature based on Doppler velocity. The compensated edge relative velocity $v_{ij}$ between points $i$ and $j$ is formulated as:
$v_{ij}=\frac{v_{i}^{dop^{{}^{\prime}}}\left\|\bm{\rho}_{i}\right\|-v_{j}^{dop^{{}^{\prime}}}\left\|\bm{\rho}_{j}\right\|}{\left\|\bm{\rho}_{i}-\bm{\rho}_{j}\right\|}.$ (4)
As Fig. 3 presents, when $\mathbf{v}_{i}^{abs}=\mathbf{v}_{j}^{abs}$, $v_{ij}$ can be further expressed as:
$v_{ij}=\mathbf{v}_{i}^{abs}\cdot\frac{(\bm{\rho}_{i}-\bm{\rho}_{j})}{\left\|\bm{\rho}_{i}-\bm{\rho}_{j}\right\|}=\mathbf{v}_{i}^{abs}\cdot\frac{\bm{\rho}_{r}}{\left\|\bm{\rho}_{r}\right\|},$ (5)
where $\bm{\rho}_{r}$ denotes the relative coordinates between $i$ and $j$.
It is evident that $v_{ij}$ remains invariant under any rotation or translation of the sensor. Therefore, if both nodes of an edge share the same absolute velocity—which holds when they belong to the same rigid object with no angular velocity—the compensated edge relative velocity $v_{ij}$ is SE(3)-invariant. Moreover, even when the two nodes have different absolute velocities, $v_{ij}$ is still SO(3)-invariant, since both $v_{i}^{dop^{\prime}}$ and $v_{j}^{dop^{\prime}}$ are individually SO(3)-invariant.
Figure 3: Illustrations of (a) node and (b) edge velocity compensation.
III-C Graph Construction
After point cloud pre-processing, the graph is constructed by treating each point as a node. A cost function $f(i,j)$ is defined to select node pairs for edge generation, preferring neighboring points with similar velocities:
$f(i,j)=\left\|\bm{\rho}_{i}-\bm{\rho}_{j}\right\|+\lambda\left\|v_{i}^{dop^{{}^{\prime}}}-v_{j}^{dop^{{}^{\prime}}}\right\|,$ (6)
where $\lambda$ is a weighting parameter. An edge is established between nodes $i$ and $j$ if $f(i,j)$ is below the threshold $F_{t}$.
Subsequently, node and edge features are specified, including the node equivariant feature $\mathbf{X}\in\mathbb{R}^{N\times 3}$, the node invariant feature $\mathbf{Y}\in\mathbb{R}^{N\times 3}$, and the edge invariant feature $\mathbf{Z}\in\mathbb{R}^{E\times 2}$, where $E$ is the total number of edges:
$\mathbf{X}$ $=\{\mathbf{X}_{i}\}=\{\bm{\rho}_{i}\},$ (7)
$\mathbf{Y}$ $=\{\mathbf{Y}_{i}\}=\{[v_{i}^{dop^{{}^{\prime}}},r_{i},d_{i}]\},$ (8)
$\mathbf{Z}$ $=\{\mathbf{Z}_{ij}\}=\{[v_{ij},\left\|\bm{\rho}_{i}-\bm{\rho}_{j}\right\|]\},j\in\mathcal{N}(i),$ (9)
where i=1,2,,Ni=1,2,...,N, $r_{i}$ denotes the RCS value of point $i$, $\mathcal{N}(i)$ is the set of neighbors of $i$ in the graph, and $d_{i}$ is the degree of node $i$, i.e., the number of edges connected to it.
The SE(3)-equivariance of $\mathbf{X}$ is straightforward, as it represents the Euclidean coordinates of the points. The SO(3)-invariance of $v_{i}^{dop^{{}^{\prime}}}$ and the SE(3)-invariance of $r_{i}$ and $d_{i}$ make the node feature $\mathbf{Y}$ invariant. Likewise, in the edge feature $\mathbf{Z}$, the compensated edge relative velocity $v_{ij}$ remains SE(3)-invariant for nodes with the same velocity, which can be reinforced by properly defining the cost function $f(i,j)$. The Euclidean distance $\left\|\bm{\rho}_{i}-\bm{\rho}_{j}\right\|$ between nodes $i$ and $j$ is also SE(3)-invariant.
III-D Network Design
Following graph construction and feature preparation, we establish a novel framework that separately processes the input equivariant and invariant features. The framework employs SE(3)-equivariant networks to handle different feature types, and estimates the relative transformation between the two frames based on the extracted final features.
III-D1 Feature Extraction
We employ two main architectures for feature extraction. The first one is an SE(3)-equivariant Dynamic Graph Convolutional Neural Network (DGCNN) [23] processing the SE(3)-equivariant node feature $\mathbf{X}$ of two consecutive point clouds and preserving its equivariance. The second one is an SE(3)-equivariant GNN (EGNN) module that, by integrating the processed node feature $\mathbf{X^{\prime}}$ with the node and edge invariant feature $\mathbf{Y}$ and $\mathbf{Z}$, outputs the final feature $\mathbf{F}$.
The first DGCNN network dynamically aggregates features from each point and its neighbors by a local graph, thereby generating richer local feature representations. Since SE(3)-equivariance is composed of translational and rotational equivariance in 3D space, we first achieve translational equivariance by translating the earlier point cloud frame to the origin of the current frame using the estimated ego-velocity and timestamp difference. This allows us to focus on rotational equivariance (i.e. SO(3)-equivariance), which is handled by a stack of LN [20] network layers in the DGCNN. LN modifies traditional MLP layers such as Linear, ReLU and Pooling layers for equivariance on semisimple Lie groups, and helps to preserve SO(3)-equivariance in our case. After three iterations, the network outputs the processed node feature $\mathbf{X^{\prime}}$, whose dimension is lifted to $N\times(C^{\prime}\times 3)$ by LN and retains SE(3)-equivariance.
The processed node feature $\mathbf{X^{\prime}}$ is then passed to the second EGNN module together with $\mathbf{Y}$ and $\mathbf{Z}$ for further graph feature aggregation. For each node $i$ and its neighbors $\mathcal{N}(i)$ connected by edges, the equivariant and invariant features are aggregated and updated as follows:
$\mathbf{m}_{ij}$ $=\phi_{I_{1}}(\exp(\gamma\left\|\mathbf{X}^{{}^{\prime}}_{i}-\mathbf{X}^{{}^{\prime}}_{j}\right\|),\mathbf{Y}_{i},\mathbf{Y}_{j},\mathbf{Z}_{ij}),$ (10)
$\mathbf{Y}_{i}$ $=\phi_{I_{2}}(\mathbf{Y}_{i},\underset{j\in\mathcal{N}(i)}{\sum}\mathbf{m}_{ij}),$ (11)
$\mathbf{X}^{{}^{\prime}}_{i}$ $=\mathbf{X}^{{}^{\prime}}_{i}+C\underset{j\in\mathcal{N}(i)}{\sum}(\mathbf{X}^{{}^{\prime}}_{i}-\mathbf{X}^{{}^{\prime}}_{j})\phi_{E}(\mathbf{m}_{ij}),$ (12)
where $\mathbf{m}_{ij}$ represents the invariant message passed from node $j$ to node $i$ under the Message Passing Neural Network (MPNN) framework [24], $\mathbf{X}^{{}^{\prime}}_{i}$ and $\mathbf{Y}_{i}$ are processed node equivariant and invariant feature of node $i$, respectively. $C$ is a constant defined as the number of elements in set $\mathcal{N}(i)$, and $\phi_{I_{1}},\phi_{I_{2}},\phi_{E}$ are MLP networks.
Differing from the original EGNN that uses relative squared distances $\left\|\mathbf{X}^{{}^{\prime}}_{i}-\mathbf{X}^{{}^{\prime}}_{j}\right\|^{2}$ for invariant feature extraction, inspired by [25], we employ a Gaussian kernel function $\exp(\gamma\left\|\mathbf{X}_{i}^{{}^{\prime}}-\mathbf{X}_{j}^{{}^{\prime}}\right\|)$ with a learnable parameter $\gamma$ to transform the SE(3)-equivariant features $\mathbf{X}$ into SE(3)-invariant representations.
After four iterations, the network outputs the final node equivariant feature set $\mathbf{X}_{F}=\{\mathbf{X}^{{}^{\prime}}_{i}\}$ and final node invariant feature set $\mathbf{Y}_{F}=\{\mathbf{Y}_{i}\}$ for i=1,2,,Ni=1,2,...,N. The final feature $\mathbf{F}$ is defined as the concatenation of $\mathbf{X}_{F}$ and $\mathbf{Y}_{F}$.
III-D2 Transformation Estimation
In this module, the $\mathbf{L}_{2}$ distances between the final features $\mathbf{F}$ of points in the two point cloud frames are computed. The $M$ pairs with the smallest feature distances are selected as keypoint correspondences. A learnable similarity metric is then defined as:
$s_{ij}=\exp(-\beta(d_{ij}-\alpha)),$ (13)
where $s_{ij}$ denotes the similarity between point $i$ in the first frame and point $j$ in the second frame, $d_{ij}$ is their feature distance, and $\alpha,\beta$ are learnable parameters that enhance the robustness of the pipeline to outliers.
Given the similarity matrix $\mathbf{S}=\{s_{ij}\}_{M\times M}$, we apply the differentiable Sinkhorn normalization algorithm [26] to obtain a soft assignment of correspondences. After $K$ iterations, the assigning weight matrix $\mathbf{W}=\{w_{ij}\}_{M\times M}$ is obtained. A weighted Singular Value Decomposition (SVD) is then used to estimate the relative transformation $(\mathbf{R},\mathbf{t})$ between the two frames.
III-D3 Loss Function
To supervise the transformation estimation, we design a novel loss function that jointly considers rotation, translation, and orientation errors. The estimated rotation $\mathbf{R}$ is first converted to Euler angles $\mathbf{E}$, and the loss $\mathcal{L}$ between $(\mathbf{E},\mathbf{t})$ and the ground truth $(\mathbf{E}_{gt},\mathbf{t}_{gt})$ consists of four terms: rotation loss $\mathcal{L}_{r}$, translation loss $\mathcal{L}_{t}$, pitch loss $\mathcal{L}_{p}$, and yaw loss $\mathcal{L}_{y}$:
$\mathcal{L}_{r}$ $=\left\|\mathbf{E}-\mathbf{E}_{gt}\right\|,$ (14)
$\mathcal{L}_{t}$ $=\left\|\mathbf{t}-\mathbf{t}_{gt}\right\|,$ (15)
$\mathcal{L}_{p}$ $=\left\|\mathbf{E}^{pitch}-\mathbf{E}_{gt}^{pitch}\right\|,$ (16)
$\mathcal{L}_{y}$ $=\left\|\mathbf{E}^{yaw}-\mathbf{E}_{gt}^{yaw}\right\|,$ (17)
where $\mathbf{E}^{pitch}$ and $\mathbf{E}_{gt}^{pitch}$ denote the pitch angles, and $\mathbf{E}^{yaw}$ and $\mathbf{E}_{gt}^{yaw}$ denote the yaw angles, in the estimated and ground truth rotations, respectively. The pitch and yaw losses are included to compensate for the relatively low elevation and azimuth resolution of 4D mmWave radar measurements.
Following [5], we introduce four learnable balancing parameters $s_{r},s_{t},s_{p},s_{y}$ to account for the different units and scales of rotation and translation, leading to the final loss:
$\mathcal{L}$ $=\mathcal{L}_{r}\exp(-s_{r})+s_{r}+\mathcal{L}_{t}\exp(-s_{t})+s_{t}$
$+\mathcal{L}_{p}\exp(-s_{p})+s_{p}+\mathcal{L}_{y}\exp(-s_{y})+s_{y}.$ (18)
IV Experiments
In this section, we present comprehensive experiments to evaluate our proposed Equi-RO framework.
IV-A Datasets
Two datasets are used for evaluation: the publicly available NTU4DRadLM dataset [6] and a self-collected dataset acquired using a vehicle-mounted 4D mmWave radar and high-precision GPS.
IV-A1 NTU4DRadLM Dataset
The NTU4DRadLM dataset is specifically designed for 4D radar-based localization and mapping. It provides accurate ground-truth odometry and includes sequences from both low-speed robot platforms and high-speed unmanned vehicles. The dataset was collected using the Oculii Eagle 4D radar, featuring an azimuth resolution of $0.5^{\circ}$, an elevation resolution of $1^{\circ}$, and an operating frequency of 12 Hz. NTU4DRadLM contains six splits. We use loop1 for training, loop2 for validation, and the remaining four splits for testing.
IV-A2 Self-Collected Dataset
To further assess generalization, we constructed a campus dataset using the Geometrical Pal R7861B 4D radar. As shown in Fig. 4, the radar was mounted on an SUV equipped with a high-precision GPS system to provide ground-truth odometry. The radar has an azimuth resolution of $1.9^{\circ}$, an elevation resolution of $3.5^{\circ}$, and operates at 15 Hz. For evaluation, models of all learning-based algorithms trained on the loop1 split of NTU4DRadLM are directly tested on this dataset to compare cross-dataset generalization.
Figure 4: Test vehicle equipped with a 4D radar for the self-collected dataset.
IV-B Experimental Setup
IV-B1 Training Details
The proposed model is trained on a single NVIDIA A100 GPU using the Adam optimizer with a weight decay of $10^{-3}$ for 50 epochs , taking approximately 6 hours. The initial learning rate is set to $10^{-4}$ and decays by a factor of 0.1 every 10 epochs. The batch size is set at 4.
IV-B2 Compared algorithms
For baseline comparison, we select several traditional point cloud registration methods, including Iterative Closest Point (ICP) [27], Normal Distributions Transform (NDT) [28], and Generalized ICP (GICP) [29]. In addition, we evaluate the State-of-the-Art (SOTA) 4D mmWave radar odometry method Adaptive Probability Distribution-GICP (APDGICP), which serves as the front end of the 4DRadarSLAM framework [2].
For learning-based methods, we include the self-supervised odometry method RaFlow [13] and the recent supervised method CAO-RONet [15].
IV-B3 Metrics
We adopt the evo library [30] to compute the relative translation error ($t_{rel}$) and relative rotation error ($r_{rel}$) for each algorithm, by comparing registered trajectories with the ground-truth.
Figure 5: Projection of adjacent point clouds onto the corresponding images, where point colors represent compensated node velocities and white circles denote selected keypoints used for matching.
Figure 6: Comparative odometry results of Equi-RO and baseline algorithms on the NTU4DRadLM dataset for the cp (a), garden (b), nyl (c), and loop3 (d) splits. Regions with large rotational motion are highlighted with magnified insets for clarity.
Figure 7: Translation error (a) and rotation error (b) profiles of Equi-RO and selected baseline algorithms over the full cp split trajectory.
Figure 8: Comparative odometry results of Equi-RO and baseline algorithms on the self-collected dataset.
Figure 9: Ablation study results of Equi-RO on the cp (a), garden (b), nyl (c), and loop3 (d) splits of the NTU4DRadLM dataset.
TABLE I: Quantitative comparison on NTU4DRadLM and Self-collected datasets ($t_{\text{rel}}:\%,r_{\text{rel}}:^{\circ}/\text{m}$)
NTU4DRadLM Self-collected
Method cp garden nyl loop3 Average 1 2 Average
$t_{\text{rel}}$ $r_{\text{rel}}$ $t_{\text{rel}}$ $r_{\text{rel}}$ $t_{\text{rel}}$ $r_{\text{rel}}$ $t_{\text{rel}}$ $r_{\text{rel}}$ $t_{\text{rel}}$ $r_{\text{rel}}$ $t_{\text{rel}}$ $r_{\text{rel}}$ $t_{\text{rel}}$ $r_{\text{rel}}$ $t_{\text{rel}}$ $r_{\text{rel}}$
ICP 6.88 0.0774 17.24 0.1997 20.05 0.0635 50.43 0.0461 23.15 0.0967 54.15 0.0698 63.89 0.3721 59.02 0.2209
NDT Failed Failed 5.46 0.0818 8.86 0.0464 13.10 0.0255 9.14 0.0512 Failed Failed Failed Failed Failed Failed
GICP 4.09 0.0576 3.27 0.0459 4.39 0.0248 4.99 0.0081 4.19 0.0341 41.23 0.0621 49.89 0.0814 45.56 0.0718
APDGICP 4.31 0.0575 3.33 0.0472 3.95 0.0210 4.81 0.0100 4.10 0.0340 10.97 0.0575 34.83 0.0961 22.90 0.0768
RaFlow 62.12 0.2362 37.07 0.3961 25.93 0.1060 60.31 0.1765 46.86 0.2287 58.43 0.0846 67.59 0.1296 63.01 0.1071
CAO-RONet 73.31 0.9404 50.42 0.5898 29.81 0.2179 78.36 0.1978 57.98 0.4865 73.31 0.0994 55.54 0.2781 64.42 0.1888
Equi-RO 3.93 0.0421 3.15 0.0380 4.31 0.0218 3.25 0.0070 3.66 0.0272 8.55 0.0524 10.23 0.0711 9.39 0.0617
IV-C Experimental Results
Table I presents a quantitative comparison of our Equi-RO against several baseline algorithms. In NTU4DRadLM dataset, our method achieves the lowest translation and rotation errors on the cp, garden, and loop3 splits, and ranks second best on the nyl split. Traditional methods such as GICP and APDGICP attain competitive results, while learning-based approaches RaFlow and CAO-RONet exhibit inferior performance. This discrepancy is likely due to their design and evaluation primarily on the earlier VoD dataset [31], leading to limited generalization on the newer NTU4DRadLM dataset. On average, our algorithm improves translation and rotation accuracy by 10.7% and 20.0%, respectively, compared to the best-performing baseline APDGICP. Figure 6 visualizes the estimated trajectories, where our approach (red lines) demonstrates superior accuracy in most splits, particularly in those involving large rotations.
As presented in Figure 5, we visualize the projection of consecutive point clouds from the loop3 split onto the corresponding images. With velocity compensation, the actual motion of dynamic points can be approximated, while the feature extraction algorithm selects keypoints that remain stationary and maintain strong continuity between frames.
Figure 7 illustrates the translation and rotation errors along the full cp split, excluding the two learning-based algorithms with large errors for clarity. While our method does not always yield the lowest translation error over the entire trajectory, it consistently achieves the best rotation accuracy. This outcome further validates the effectiveness of our invariant velocity preprocessing, SE(3)-equivariant network design, and adaptive loss function in robustly encoding and estimating rotational motion.
The comparative results on our self-collected dataset are also summarized in Table I and visualized in Figure 8. Due to the lower angular resolution of the employed 4D radar compared to that in the NTU4DRadLM dataset, all methods experience performance degradation. Nonetheless, our approach still attains the best translation and rotation accuracy across both splits, despite being trained solely on the NTU4DRadLM dataset, which demonstrates its strong generalization capability.
IV-D Ablation Study
To further validate the effectiveness of our proposed algorithm, we conduct an ablation study on the NTU4DRadLM dataset by creating three variant versions of Equi-RO:
  • Equi-RO with No Invariant Velocity Compensation (NIV): In this variant, node velocities are taken directly from the original Doppler values $v_{i}^{dop}$, and edge velocities are computed as $v_{i}^{dop}-v_{j}^{dop}$, rather than using the compensated counterparts $v_{i}^{dop^{\prime}}$ and $v_{ij}$.
  • Equi-RO with No Equivariant Network Design (NEN): Here, the LN-based layers are replaced by conventional Linear, ReLU, and Pooling layers, and the EGNN is substituted with a standard MPNN, removing the equivariance property from the network.
  • Equi-RO with No Adaptive Loss (NAL): In this case, the pitch loss $\mathcal{L}_{p}$ and yaw loss $\mathcal{L}_{y}$ are omitted, and fixed weights ($s_{r}=-8.0$, $s_{t}=-3.0$) are assigned to the rotation and translation losses instead of using learnable balancing parameters.
The quantitative results of this ablation study are summarized in Table II.
TABLE II: Ablation study results on the NTU4DRadLM dataset
Method cp garden nyl loop3
$t_{\text{rel}}$ $r_{\text{rel}}$ $t_{\text{rel}}$ $r_{\text{rel}}$ $t_{\text{rel}}$ $r_{\text{rel}}$ $t_{\text{rel}}$ $r_{\text{rel}}$
NIV 5.16 0.0542 4.92 0.0601 8.12 0.0432 4.31 0.0088
NEN 9.60 0.1688 13.27 0.1960 8.97 0.0376 7.97 0.0101
NAL 5.32 0.0498 4.76 0.0744 9.11 0.0504 5.24 0.0222
Equi-RO 3.93 0.0421 3.15 0.0380 4.31 0.0218 3.25 0.0070
We also visualize the trajectories generated by each ablated variant across different splits in Fig. 9. These trajectories demonstrate that the full Equi-RO model consistently achieves the best performance. All three ablated variants show notable performance degradation, particularly in scenarios involving large rotations. The NEN variant, which discards equivariant network design, exhibits the worst results across most splits, underscoring the critical importance of equivariant architecture in our framework.
IV-E Limitations
Although Equi-RO demonstrates superior accuracy and robustness, its performance is affected by the relatively low resolution of radars, as shown in the self-collected dataset. Additionally, while our approach meets general real-time requirements with a processing speed of 76.75 ms per frame on a single NVIDIA A100 GPU, further algorithmic optimization is necessary for deployment on on-board platforms with stricter latency budgets.
V Conclusion
In this paper, we propose a novel 4D mmWave radar odometry algorithm based on equivariant networks. Our method effectively handles noisy, sparse radar point clouds while preserving geometric consistency under large rotations. Extensive experiments on NTU4DRadLM and the self-collected dataset validate its effectiveness and generalizability. This work highlights the potential of equivariant network design for radar-based perception and provides a new perspective for point cloud processing in robotics and autonomous systems. Future work will extend the framework to multi-modal fusion with IMUs and cameras and explore its application to other 3D point clouds with physical attributes.

### v2

@@
 A graph-based architecture enhances feature aggregation in sparse radar data, improving inter-frame correspondence.
-Experiments on the open-source dataset and self-collected dataset show Equi-RO outperforms state-of-the-art algorithms in accuracy and robustness.
-Overall, our method achieves 10.7% and 20.0% relative improvements in translation and rotation accuracy, respectively, compared to the best baseline on the open-source dataset.
+Experiments on an open-source dataset and a self-collected dataset show Equi-RO outperforms state-of-the-art algorithms in accuracy and robustness.
+Overall, our method achieves 10.7% and 13.4% relative improvements in translation and rotation accuracy, respectively, compared to the best baseline on the open-source dataset.
 I Introduction
-Odometry is indispensable for autonomous vehicles and robots, providing accurate support for downstream perception, planning, and control.
-When Global Positioning System (GPS) signals are degraded, such as under bridges, in tunnels, or during severe rain or snow, it becomes crucial to obtain relative localization using sensors such as LiDARs and cameras.
-However, the performance of LiDARs and cameras degrades significantly under extreme weather conditions, limiting their applicability for odometry estimation in such scenarios.
-Recently, 4D millimeter-wave (mmWave) radar—characterized by compact size, cost efficiency, all-weather adaptability, velocity-measuring capability, and long detection range—has been increasingly regarded as a promising solution for robust perception and localization in autonomous driving, particularly under adverse weather conditions [1].
-Current 4D mmWave radar odometry algorithms are mostly adapted from traditional LiDAR odometry methods.
-Some exploit unique properties of 4D mmWave radar point clouds like Doppler velocity and Radar Cross Section (RCS) to perform ego-velocity estimation and point cloud registration [2, 3].
-Regarding learning-based approaches, a few pioneering works have adapted Convolutional Neural Networks (CNNs) and Recurrent Neural Networks (RNNs) for stand-alone 4D mmWave radar [4] and radar-camera fusion [5] odometry.
+Odometry is indispensable for autonomous vehicles and robots, supporting downstream perception, planning, and control.
+When Global Positioning System (GPS) signals are degraded, localization relies primarily on odometry from onboard sensors such as LiDARs and cameras.
+However, these sensors often struggle in extreme weather, whereas 4D millimeter-wave (mmWave) radar (4D radar) provides compact, cost-efficient, all-weather, velocity-measuring, and long-range sensing, making it a promising alternative for robust perception and localization under adverse conditions [12].
+Current 4D radar odometry mostly adapts traditional LiDAR methods and exploits Doppler velocity and Radar Cross Section (RCS) for ego-velocity estimation and point cloud registration [35, 38], while a few learning-based works employ convolutional and recurrent neural networks for 4D radar [21] and radar-camera fusion [39] odometry.
 Figure 1: Our Equi-RO method incorporates Doppler velocity compensation and graph construction to preserve noise-resilient and rotation-robust features.
-The derived equivariant and invariant features are fed into an equivariant Graph Neural Network (GNN)-based framework to obtain consistent and generalizable point cloud registration, yielding more accurate odometry results, especially in large rotations.
-Trajectory comparison on a partial segment of the loop3 split from the NTU4DRadLM dataset[6], with initialization aligned to the ground truth at the segment’s start, against [2].
-In learning-based odometry, accurate feature extraction and matching are critical steps.
+The derived equivariant and invariant features are fed into an equivariant Graph Neural Network (GNN)-based framework for consistent and generalizable point cloud registration, yielding more accurate odometry results, especially in large rotations.
+Trajectory comparison on a partial segment of the loop3 split from the NTU4DRadLM dataset [34], where both methods are initialized to the ground truth at the segment’s start, comparing against [35].
+In learning-based odometry, feature extraction and matching are critical.
 However, features in adjacent frames may become highly inconsistent when the ego-vehicle moves rapidly or turns sharply.
 Equivariant neural networks have emerged as a promising approach to address this issue [5, 37].
-They learn features that transform predictably under specific geometric transformations, such as 3D translations and rotations represented by rigid-body motion group SE(3), thereby enhancing feature matching performance.
-The geometric symmetry can improve the generalizability and interpretability of learning-based odometry systems.
-The 4D mmWave radar provides additional physical attributes compared to LiDAR point clouds.
-For example, it measures the RCS of each point, which can be regarded as SE(3)-invariant since it reflects the strength of the returned signal and is determined mainly by the target’s size, shape, and material rather than the sensor’s pose.
-In addition, the Doppler effect enables estimation of the radial relative velocity of detected targets.
-However, Doppler velocity is neither equivariant nor invariant under SE(3) transformations, as it depends on both the sensor’s motion and the target’s motion, which limits the applicability of existing equivariant networks to 4D mmWave radar odometry.
-Moreover, 4D mmWave radar point clouds are inherently noisier, sparser, and more irregular than those captured by scanning LiDARs.
-Consequently, effectively aggregating point-wise features with those of their spatial neighbors remains a significant challenge for learning-based 4D mmWave radar odometry.
-In this paper, we propose Equi-RO, a novel framework for 4D mmWave radar odometry built upon equivariant neural networks.
-Compared with existing 4D mmWave radar odometry algorithms, we design a pre-processing pipeline to derive invariant Doppler features and construct a graph representation to mitigate noise and sparsity.
-The equivariant and invariant features are then processed through separate networks to estimate the relative transformation.
+They learn features that transform predictably under, e.g., rigid body transformation group SE(3) (equivariance property), or remain fixed under such transformations (invariance property).
+In 4D radar sensing, the Doppler effect enables estimation of targets’ radial relative velocity.
+However, Doppler velocity is neither equivariant nor invariant under the sensor’s transformations, as it depends on both the sensor’s and target’s motion, limiting the applicability of equivariant networks to 4D radar odometry.
+Moreover, 4D radar point clouds, with typically only thousands of points per frame, are noisier, sparser, and more irregular than LiDAR point clouds.
+Consequently, effectively aggregating point-wise features from spatial neighbors remains a challenge for learning-based 4D radar odometry.
+In this paper, we propose Equi-RO, a novel 4D radar odometry framework built upon equivariant networks (Fig.
+1).
+We design a pre-processing pipeline to derive invariant Doppler features and construct a graph from the point cloud to mitigate noise and sparsity.
+Compared with raw point sets or voxel grids, a graph representation explicitly connects spatially and kinematically consistent neighbors, which can suppress isolated noise and stabilize feature extraction for 4D radar point clouds.
 The invariant velocity features and equivariant network design enable our method to remain robust in challenging real-world driving scenarios, where existing methods often degrade.
-Leveraging both a public dataset [6] and a self-collected dataset, we demonstrate consistent improvements over state-of-the-art 4D mmWave radar odometry approaches.
+Leveraging both a public dataset [34] and a self-collected dataset, we demonstrate consistent improvements over state-of-the-art 4D radar odometry approaches.
 In summary, our main contributions are as follows:
   1. 1.
-    We develop a novel 4D mmWave radar odometry algorithm based on equivariant neural networks, capable of producing accurate and robust odometry estimates under large rotations and fast motions, as commonly encountered in sharp turns and challenging driving maneuvers.
+    We develop a novel 4D radar odometry algorithm based on equivariant neural networks, providing accurate and robust odometry estimates under large rotations, as commonly encountered in sharp turns and challenging driving maneuvers.
   2. 2.
     We design a dedicated pre-processing pipeline for Doppler velocity that derives invariant features, thereby stabilizing motion-related features.
-And we design a framework that separately extracts equivariant and invariant features to preserve geometric consistency, making it applicable to 4D mmWave radar point clouds with physical attributes.
+We also design a framework that separately extracts equivariant and invariant features to preserve geometric consistency, making it applicable to 4D radar point clouds with physical attributes.
   3. 3.
     We conduct extensive experiments on both an open-source dataset and a self-collected dataset, along with ablation studies, to validate the effectiveness of the proposed method across diverse scenarios.
-  4. 4.
-    The source code will be released after receiving the final decision.
+The project is available at: https://github.com/hanzy21/Equi-RO.
 II Related Works
-In this section, we briefly review existing algorithms for 4D mmWave radar odometry, followed by a discussion of equivariant neural networks.
-II-A 4D mmWave Radar Odometry
-Existing 4D mmWave radar odometry approaches can be broadly categorized into traditional and learning-based methods.
-Traditional odometry methods, which do not involve neural networks, primarily address two challenges: (1) the effective utilization of Doppler velocity and RCS information, and (2) the handling of noise and sparsity in radar point clouds.
-For the first aspect, Doppler velocity is commonly employed for ego-velocity estimation [9] and dynamic object removal [3], while RCS is often used for point cloud filtering [10] and weighted point matching [11].
-In the latter case, some works have proposed to use scan-to-submap registration [3] and adaptive probability-based registration [2, 12] to handle the noise and sparsity.
-Learning-based approaches replace traditional point cloud registration and pose regression with deep neural networks.
-As an early pioneer, Lu et al.
-[4] employed CNNs to extract features from radar point clouds.
-RaFlow [13] formulates odometry estimation by scene flow prediction.
-4DRO-Net [14] and 4DRVO-Net [5] integrate velocity-aware attention cost volumes into a coarse-to-fine optimization framework to iteratively estimate poses.
-CAO-RONet [15] incorporates local completion and context-aware association to match noisy points and suppress outliers.
-To the best of our knowledge, although GNNs can explicitly aggregate features from spatially neighboring points, they have not yet been applied to 4D mmWave radar odometry.
-II-B Equivariant Neural Networks
-Equivariant neural networks are designed to produce predictable changes in their outputs when specific transformations are applied to the inputs.
-This property enhances both the generalizability and interpretability of the models.
-Since Cohen and Welling first introduced the concept of equivariance into neural networks [7], various equivariant designs have been proposed for GNNs [16, 17].
-For example, Satorras et al.
-[17] propose Equivariant Graph Neural Network (EGNN), which maintains equivariance to the Euclidean group E(n)—including translations, rotations, and reflections—by incorporating the relative squared distances between coordinates in the edge feature aggregation process.
-Significant efforts have also been devoted to designing equivariant architectures for point cloud analysis.
-Tensor Field Network (TFN) [18] employs spherical harmonics to construct learnable weight kernels, preserving equivariance to SE(3) for point cloud data.
-Vector Neurons (VN) [19] and Lie Neurons (LN) [20] extend traditional Multi-Layer Perceptrons (MLPs) to operate on vector features, achieving equivariance to 3D rotation group SO(3) and semisimple Lie groups, respectively.
+In this section, we briefly review existing algorithms for 3D point cloud odometry and 4D radar odometry, followed by a discussion of equivariant neural networks.
+II-A 3D Point Cloud Odometry
+3D point cloud odometry for LiDAR and RGB-D cameras can be categorized into traditional and learning-based methods.
+Traditional methods mainly utilize geometric registration, such as Iterative Closest Point (ICP) [1] and Normal Distributions Transform (NDT) [2].
+Subsequent Generalized-ICP (GICP) [27] and KISS-ICP [31] further improve robustness through local surface modeling.
+Learning-based methods such as FCGF [4] and DeepVCP [23] replace handcrafted features with learned correspondence models on dense point clouds.
+Despite their success on dense point clouds, these methods’ heavy reliance on dense and smooth geometry limits their applicability to sparse and noisy data from 4D radars.
+II-B 4D Radar Odometry
+Several core ideas of 3D point cloud odometry including geometric registration and local feature extraction are transferable to 4D radar odometry.
+However, leveraging radar-specific measurements (Doppler and RCS) and handling the sparsity and noise of radar point clouds remain open challenges.
+These issues have led to recent works on 4D radar odometry.
+For measurements utilization, Doppler velocity is commonly employed for ego-velocity estimation [8, 32], dynamic object removal [38, 14] and velocity-aware attention module design [22, 39], while RCS is often used for point cloud filtering [15, 13] and weighted point matching [16] since it indicates the intensity of the point.
+As for handling the sparsity and noise of radar point clouds, classical 4D radar odometry uses scan-to-submap [38, 17], adaptive probability-based [35, 18], and voxel-based [28] registration to improve the robustness.
+For learning-based methods, RaFlow [7] predicts scene flow, while CAO-RONet [19] incorporates local completion to match noisy points.
+To the best of our knowledge, although GNNs can explicitly aggregate features from spatially neighboring points, they have not yet been applied to 4D radar odometry.
+II-C Equivariant Neural Networks
+Equivariant neural networks produce predictable outputs under specific input transformations, enhancing generalizability and interpretability.
+Since Cohen and Welling introduced group equivariance into neural networks [5], various equivariant designs have been proposed for GNNs [9, 26].
+For example, Equivariant Graph Neural Network (EGNN) [26] maintains equivariance to the Euclidean group E(n), which includes translations, rotations, and reflections, by incorporating the relative distances in the edge feature aggregation.
+For point cloud analysis, Tensor Field Network [30] achieves SE(3)-equivariance via spherical harmonics, while Vector Neurons (VN) [6] and Lie Neurons (LN) [20] extend traditional Multi-Layer Perceptrons (MLPs) to vector features, achieving equivariance to 3D rotation group SO(3) and semi-simple Lie groups, respectively.
 EPN [3] and E2PN [37] further introduce SE(3)-equivariant convolutional architectures for 3D point clouds.
-Despite these advances, the application of equivariant neural networks to 4D mmWave radar point clouds remains largely unexplored.
+However, equivariant models remain largely underexplored for 4D radar point clouds.
 III Methodology
 This section provides a detailed description of our proposed Equi-RO algorithm.
-We first present an overview of the approach in Sec.
-III-A, followed by the pre-processing pipeline for 4D mmWave radar point clouds in Sec.
-III-B.
-Sec.
-III-C introduces the graph construction process and the definition of node and edge features, while Sec.
-III-D outlines the overall network architecture of Equi-RO.
 III-A Overview
-The overall pipeline of Equi-RO is illustrated in Fig.
-2.
-Two consecutive frames of 4D mmWave radar point clouds are first pre-processed to estimate the ego-vehicle’s velocity and to generate invariant compensated node and edge velocities based on Doppler velocity.
-The processed point clouds are represented as graphs, with carefully defined node and edge features.
-An SE(3)-equivariant network based on LN [20] is employed to extract and aggregate node-level equivariant features, while both node- and edge-level invariant features are fed into an EGNN-based [17] module.
-The combined features are ultimately used to estimate the relative transformation $(R,t)$ between the two input frames, achieving robust registration even under large rotations.
+Fig.
+2 shows the pipeline of Equi-RO.
+Standard 4D radar odometry faces two primary challenges: the sparsity of radar point clouds and the ego-motion dependent nature of Doppler velocities.
+To address sparsity and noise, we structure the point cloud as a graph rather than isolated points.
+This allows us to explicitly define edge features that capture local relative patterns, aggregating information from neighbors to improve robustness.
+To address the motion coupling of Doppler velocity, we mathematically derive invariant compensated velocities to decouple the object’s motion from the ego-vehicle’s rotation.
+Based on these pre-processed inputs, we extract equivariant geometric structures (e.g., 3D coordinates that transform consistently with the sensor) and invariant physical features (e.g., RCS and our derived Doppler velocity that remain unchanged under the sensor’s rotation or translation), and process these features with different networks.
+The equivariant features enable the network to reason about relative motion during scan matching, while the invariant features remain consistent across different sensor orientations, providing pose-independent cues.
+Specifically, an SO(3)-equivariant network based on LN [20] is employed to extract node-level equivariant features, while both node- and edge-level invariant features are fed into an EGNN-based [26] module.
+The output features are used to estimate the relative transformation $(\bm{R},\bm{t})$ between the two input frames, achieving robust registration even under large rotations.
 Figure 2: Overview of the Equi-RO algorithm.
-Doppler velocity is used to derive invariant node and edge features, which are combined with node and edge features within a unified graph-based network framework to robustly estimate the relative transformation $(R,t)$ between two consecutive radar frames.
+Doppler velocity is used to derive invariant node and edge features, which are combined with node and edge features within a unified graph-based network framework to robustly estimate the relative transformation $(\bm{R},\bm{t})$ between two consecutive radar frames.
 III-B Point Cloud Pre-processing
-As mentioned above, 4D mmWave radar can directly measure the radial relative velocity of detected targets via the Doppler effect.
-In this section, we fully exploit the Doppler velocity to estimate the ego-vehicle’s velocity and to derive compensated node and edge velocities for subsequent use in the GNN-based equivariant framework.
+As mentioned above, 4D radar can directly measure the radial relative velocity of targets via the Doppler effect.
+In this section, we fully exploit the Doppler velocity to estimate the ego-velocity and to derive compensated node and edge velocities for subsequent GNN-based equivariant framework.
 III-B1 Ego-velocity Estimation
 We apply an iterative reweighted least squares algorithm [33] to estimate the ego-velocity $\mathbf{v}_{ego}$.
-The Doppler velocity $v_{i}^{dop}$ of each point $i$ is affected by the coordinate $\bm{\rho}_{i}$ of this point in the sensor frame.
-In this work, Doppler velocity is defined as positive for approaching targets and negative for receding ones.
-For a point with absolute velocity $\mathbf{v}_{i}^{abs}$, its Doppler velocity $v_{i}^{dop}$ can be represented as:
+For a point $i$ with absolute velocity $\mathbf{v}_{i}^{abs}$, its Doppler velocity $v_{i}^{dop}$ can be represented as:
 $v_{i}^{dop}=\frac{\bm{\rho}_{i}^{T}}{\left\|\bm{\rho}_{i}\right\|}\cdot(\mathbf{v}_{ego}-\mathbf{v}_{i}^{abs}),$ (1)
-where $\bm{\rho}_{i}^{T}$ is the transpose of $\bm{\rho}_{i}$.
-Assuming that most points in the scene are stationary, and assigning each point $i$ a weight $\omega_{i}$ indicating its likelihood of being stationary, we formulate the following weighted least squares problem for a point cloud with $N$ points:
-$\underset{\mathbf{v}_{ego}}{\min}\sum_{i=1}^{N}\omega_{i}\left\|v_{i}^{dop}-\frac{\bm{\rho}_{i}^{T}}{\left\|\bm{\rho}_{i}\right\|}\cdot\mathbf{v}_{ego}\right\|.$ (2)
-Here, the weights are iteratively updated as $\omega_{i}=1/(v_{r}+\epsilon)$, where $v_{r}=\left\|v_{i}^{dop}-\frac{\bm{\rho}_{i}^{T}}{\left\|\bm{\rho}_{i}\right\|}\cdot\mathbf{v}_{ego}\right\|$ and $\epsilon=1\times 10^{-5}$.
-Upon convergence, $\mathbf{v}_{ego}$ provides a robust estimate of the ego-vehicle’s velocity.
+where $\bm{\rho}_{i}$ is the coordinate of this point.
+We define the Doppler residual $v_{i}^{r}=\left\|v_{i}^{dop}-\frac{\bm{\rho}_{i}^{T}}{\left\|\bm{\rho}_{i}\right\|}\cdot\mathbf{v}_{ego}\right\|$, which should be zero for stationary points.
+To robustly suppress moving points, we can formulate the following weighted least squares problem for a point cloud with $N$ points:
+$\underset{\mathbf{v}_{ego}}{\min}\sum_{i=1}^{N}\omega_{i}{v_{i}^{r}}^{2}.$ (2)
+The weight $\omega_{i}$ indicates the likelihood of a point being stationary and is iteratively updated as $\omega_{i}=1/(v_{i}^{r}+\epsilon)$, with $\epsilon=1\times 10^{-5}$.
+Upon convergence, $\mathbf{v}_{ego}$ yields a robust ego-velocity estimate.
 III-B2 Node Velocity Compensation
-Once $\mathbf{v}_{ego}$ is estimated, the compensated node velocity $v_{i}^{dop^{\prime}}$ for each point $i$ can be computed as:
-$v_{i}^{dop^{{}^{\prime}}}=v_{i}^{dop}-\frac{\bm{\rho}_{i}^{T}}{\left\|\bm{\rho}_{i}\right\|}\cdot\mathbf{v}_{ego}.$ (3)
+Once $\mathbf{v}_{ego}$ is estimated, the compensated node velocity $v_{i}^{dop^{\prime}}$ for each point $i$ is:
+$v_{i}^{dop^{{}^{\prime}}}=v_{i}^{dop}-\frac{\bm{\rho}_{i}^{T}}{\left\|\bm{\rho}_{i}\right\|}\cdot\mathbf{v}_{ego}=-\frac{\bm{\rho}_{i}^{T}}{\left\|\bm{\rho}_{i}\right\|}\cdot\mathbf{v}_{i}^{abs}.$ (3)
 As illustrated in Fig.
-3, $v_{i}^{dop^{\prime}}$ corresponds to $\frac{\bm{\rho}_{i}^{T}}{\left\|\bm{\rho}_{i}\right\|}\cdot\mathbf{v}_{i}^{abs}$, i.e., the radial component of the absolute velocity $\mathbf{v}_{i}^{abs}$ of point $i$.
-Notably, $v_{i}^{dop^{\prime}}$ is invariant under SO(3) transformations of the radar, providing a stable feature representation for point cloud registration under large rotational variations.
-Since $\mathbf{v}_{ego}$ has already been obtained to enhance relative translation estimation, incorporating the SO(3)-invariant $v_{i}^{dop^{\prime}}$ further strengthens the robustness of feature matching, thereby improving rotational alignment accuracy.
+3 (left), $v_{i}^{dop^{\prime}}$ is the radial component of the absolute velocity $\mathbf{v}_{i}^{abs}$ of point $i$.
+If the radar coordinate frame is rotated by a matrix $\bm{R}\in SO(3)$ with $\bm{R}^{T}\bm{R}=\bm{I}$, $v_{i}^{dop^{\prime}}$ remains the same as:
+$v_{i}^{dop^{{}^{\prime\prime}}}$ $=-\frac{(\bm{R\rho}_{i})^{T}}{\left\|\bm{R\rho}_{i}\right\|}\cdot\bm{R}\mathbf{v}_{i}^{abs}=-\frac{\bm{\rho}_{i}^{T}\bm{R}^{T}\bm{R}\mathbf{v}_{i}^{abs}}{\left\|\bm{\rho}_{i}\right\|}$
+$=-\frac{\bm{\rho}_{i}^{T}}{\left\|\bm{\rho}_{i}\right\|}\cdot\mathbf{v}_{i}^{abs}=v_{i}^{dop^{{}^{\prime}}}.\vskip-14.22636pt$ (4)
+Therefore, $v_{i}^{dop^{\prime}}$ is invariant under SO(3) transformations of the radar, providing a stable feature representation for point cloud registration under large rotational variations.
 III-B3 Edge Velocity Compensation
@@
 $v_{ij}=\frac{v_{i}^{dop^{{}^{\prime}}}\left\|\bm{\rho}_{i}\right\|-v_{j}^{dop^{{}^{\prime}}}\left\|\bm{\rho}_{j}\right\|}{\left\|\bm{\rho}_{i}-\bm{\rho}_{j}\right\|}.$ (5)
-As Fig. 3 presents, when $\mathbf{v}_{i}^{abs}=\mathbf{v}_{j}^{abs}$, $v_{ij}$ can be further expressed as:
-$v_{ij}=\mathbf{v}_{i}^{abs}\cdot\frac{(\bm{\rho}_{i}-\bm{\rho}_{j})}{\left\|\bm{\rho}_{i}-\bm{\rho}_{j}\right\|}=\mathbf{v}_{i}^{abs}\cdot\frac{\bm{\rho}_{r}}{\left\|\bm{\rho}_{r}\right\|},$ (5)
-where $\bm{\rho}_{r}$ denotes the relative coordinates between $i$ and $j$.
-It is evident that $v_{ij}$ remains invariant under any rotation or translation of the sensor.
-Therefore, if both nodes of an edge share the same absolute velocity—which holds when they belong to the same rigid object with no angular velocity—the compensated edge relative velocity $v_{ij}$ is SE(3)-invariant.
-Moreover, even when the two nodes have different absolute velocities, $v_{ij}$ is still SO(3)-invariant, since both $v_{i}^{dop^{\prime}}$ and $v_{j}^{dop^{\prime}}$ are individually SO(3)-invariant.
-Figure 3: Illustrations of (a) node and (b) edge velocity compensation.
+Since both $v_{i}^{dop^{\prime}}$ and $v_{j}^{dop^{\prime}}$ are individually SO(3)-invariant, and rotations preserve vector norms and pairwise distances, it is evident that $v_{ij}$ is also SO(3)-invariant.
+Moreover, as Fig.
+3 (right) presents, if both nodes of an edge have identical absolute velocities, i.e., $\mathbf{v}_{i}^{abs}=\mathbf{v}_{j}^{abs}$, we can substitute Eq.
+(3) into Eq.
+(5), $v_{ij}$ as:
+$v_{ij}=\frac{(\bm{\rho}_{j}^{T}-\bm{\rho}_{i}^{T})}{\left\|\bm{\rho}_{i}-\bm{\rho}_{j}\right\|}\cdot\mathbf{v}_{i}^{abs}=\frac{\bm{\rho}_{r}^{T}}{\left\|\bm{\rho}_{r}\right\|}\cdot\mathbf{v}_{i}^{abs},$ (6)
+where $\bm{\rho}_{r}$ denotes the relative coordinates from $i$ to $j$.
+If the radar undergoes a rigid-body transformation $(\bm{R},\bm{t})\in SE(3)$, the translation $\bm{t}$ does not change $\mathbf{v}_{i}^{abs}$ and $\bm{\rho_{r}}$.
+So we have:
+$v_{ij}^{{}^{\prime}}$ $=\frac{(\bm{R\rho}_{r})^{T}}{\left\|\bm{R\rho}_{r}\right\|}\cdot(\bm{R}\mathbf{v}_{i}^{abs})=\frac{\bm{\rho}_{r}^{T}\bm{R}^{T}\bm{R}\mathbf{v}_{i}^{abs}}{\left\|\bm{\rho}_{r}\right\|}$ (7)
+$=\frac{\bm{\rho}_{r}^{T}}{\left\|\bm{\rho}_{r}\right\|}\cdot\mathbf{v}_{i}^{abs}=v_{ij}.$
+Hence, the edge compensated velocity $v_{ij}$ is SE(3)-invariant when its two nodes share the same velocity, which holds when they belong to a rigid object with no angular velocity.
+Figure 3: Illustrations of (left) node and (right) edge velocity compensation.
+The yellow arrow in the left figure indicates the compensated node velocity, which remains SO(3)-invariant.
+The yellow arrow in the right figure is the compensated edge velocity when the two nodes share the same absolute velocity, which is SE(3)-invariant.
 III-C Graph Construction
-After point cloud pre-processing, the graph is constructed by treating each point as a node.
+After pre-processing, the graph is constructed for each frame of point cloud by treating each point as a node.
 A cost function $f(i,j)$ is defined to select node pairs for edge generation, preferring neighboring points with similar velocities:
 $f(i,j)=\left\|\bm{\rho}_{i}-\bm{\rho}_{j}\right\|+\lambda\left\|v_{i}^{dop^{{}^{\prime}}}-v_{j}^{dop^{{}^{\prime}}}\right\|,$ (8)
-where $\lambda$ is a weighting parameter. An edge is established between nodes $i$ and $j$ if $f(i,j)$ is below the threshold $F_{t}$.
-Subsequently, node and edge features are specified, including the node equivariant feature $\mathbf{X}\in\mathbb{R}^{N\times 3}$, the node invariant feature $\mathbf{Y}\in\mathbb{R}^{N\times 3}$, and the edge invariant feature $\mathbf{Z}\in\mathbb{R}^{E\times 2}$, where $E$ is the total number of edges:
-$\mathbf{X}$ $=\{\mathbf{X}_{i}\}=\{\bm{\rho}_{i}\},$ (7)
-$\mathbf{Y}$ $=\{\mathbf{Y}_{i}\}=\{[v_{i}^{dop^{{}^{\prime}}},r_{i},d_{i}]\},$ (8)
-$\mathbf{Z}$ $=\{\mathbf{Z}_{ij}\}=\{[v_{ij},\left\|\bm{\rho}_{i}-\bm{\rho}_{j}\right\|]\},j\in\mathcal{N}(i),$ (9)
-where i=1,2,,Ni=1,2,...,N, $r_{i}$ denotes the RCS value of point $i$, $\mathcal{N}(i)$ is the set of neighbors of $i$ in the graph, and $d_{i}$ is the degree of node $i$, i.e., the number of edges connected to it.
-The SE(3)-equivariance of $\mathbf{X}$ is straightforward, as it represents the Euclidean coordinates of the points.
-The SO(3)-invariance of $v_{i}^{dop^{{}^{\prime}}}$ and the SE(3)-invariance of $r_{i}$ and $d_{i}$ make the node feature $\mathbf{Y}$ invariant.
-Likewise, in the edge feature $\mathbf{Z}$, the compensated edge relative velocity $v_{ij}$ remains SE(3)-invariant for nodes with the same velocity, which can be reinforced by properly defining the cost function $f(i,j)$.
-The Euclidean distance $\left\|\bm{\rho}_{i}-\bm{\rho}_{j}\right\|$ between nodes $i$ and $j$ is also SE(3)-invariant.
+where $\lambda$ is a weighting parameter. Edges are established between node $i$ and its $k$ nearest neighbors in the feature space.
+Subsequently, to effectively fuse spatial and physical information, we organize the input into three distinct sets based on their transformation properties:
+Geometric equivariant node feature set $\mathbf{X}\in\mathbb{R}^{N\times 3}$: This set contains Euclidean coordinates $\bm{\rho}_{i}$ of each node $i$ which reflect the geometric information and are inherently SE(3)-equivariant: $\mathbf{X}=\{\mathbf{X}_{i}\}=\{\bm{\rho}_{i}\}$.
+Physical invariant node feature set $\mathbf{Y}\in\mathbb{R}^{N\times 3}$: This set represents the intrinsic properties of each node $i$ that remain stable under transformation, including the SO(3)-invariant compensated node velocity $v_{i}^{dop^{{}^{\prime}}}$, and SE(3)-invariant RCS value $r_{i}$ and the degree $d_{i}$, which is the number of edges connected to the node $i$: $\mathbf{Y}=\{\mathbf{Y}_{i}\}=\{[v_{i}^{dop^{{}^{\prime}}},r_{i},d_{i}]\}$.
+Relative invariant edge feature set $\mathbf{Z}\in\mathbb{R}^{E\times 2}$ ($E$ is the total number of edges): To capture local relative information between nodes $i$ and $j$, the edge features consist of the compensated edge relative velocity $v_{ij}$, which remains SO(3)-invariant and further satisfies SE(3)-invariance for nodes with the same velocity, and the SE(3)-invariant Euclidean distance $\left\|\bm{\rho}_{i}-\bm{\rho}_{j}\right\|$: $\mathbf{Z}=\{\mathbf{Z}_{ij}\}=\{[v_{ij},\left\|\bm{\rho}_{i}-\bm{\rho}_{j}\right\|]\},j\in\mathcal{N}(i)$, where $\mathcal{N}(i)$ is the set of neighbors of $i$ in the graph.
+These constructed feature sets ($\mathbf{X},\mathbf{Y},\mathbf{Z}$), and the graph structure $\mathbf{G}$, are built independently for both the source and target frames (denoted by subscripts $s$ and $t$ in Fig.
+2) to serve as paired inputs for the subsequent network.
 III-D Network Design
-Following graph construction and feature preparation, we establish a novel framework that separately processes the input equivariant and invariant features.
-The framework employs SE(3)-equivariant networks to handle different feature types, and estimates the relative transformation between the two frames based on the extracted final features.
+Following graph construction and feature preparation, we establish a novel framework that processes the input equivariant and invariant features for both the source and target point clouds independently.
+The framework employs equivariant networks to handle different feature types and progressively integrate geometric structure with physical properties, yielding the respective final per-point feature set $\mathbf{F}_{s}$ and $\mathbf{F}_{t}$ for estimating the relative transformation between the two frames.
 III-D1 Feature Extraction
-We employ two main architectures for feature extraction.
-The first one is an SE(3)-equivariant Dynamic Graph Convolutional Neural Network (DGCNN) [23] processing the SE(3)-equivariant node feature $\mathbf{X}$ of two consecutive point clouds and preserving its equivariance.
-The second one is an SE(3)-equivariant GNN (EGNN) module that, by integrating the processed node feature $\mathbf{X^{\prime}}$ with the node and edge invariant feature $\mathbf{Y}$ and $\mathbf{Z}$, outputs the final feature $\mathbf{F}$.
-The first DGCNN network dynamically aggregates features from each point and its neighbors by a local graph, thereby generating richer local feature representations.
-Since SE(3)-equivariance is composed of translational and rotational equivariance in 3D space, we first achieve translational equivariance by translating the earlier point cloud frame to the origin of the current frame using the estimated ego-velocity and timestamp difference.
-This allows us to focus on rotational equivariance (i.e.
-SO(3)-equivariance), which is handled by a stack of LN [20] network layers in the DGCNN.
-LN modifies traditional MLP layers such as Linear, ReLU and Pooling layers for equivariance on semisimple Lie groups, and helps to preserve SO(3)-equivariance in our case.
-After three iterations, the network outputs the processed node feature $\mathbf{X^{\prime}}$, whose dimension is lifted to $N\times(C^{\prime}\times 3)$ by LN and retains SE(3)-equivariance.
-The processed node feature $\mathbf{X^{\prime}}$ is then passed to the second EGNN module together with $\mathbf{Y}$ and $\mathbf{Z}$ for further graph feature aggregation.
-For each node $i$ and its neighbors $\mathcal{N}(i)$ connected by edges, the equivariant and invariant features are aggregated and updated as follows:
+We employ a two-stage cascaded architecture for feature extraction of each frame.
+An SO(3)-equivariant Dynamic Graph Convolutional Neural Network (DGCNN) [25] first processes equivariant node feature set $\mathbf{X}$ of each frame of point clouds and produces SO(3)-equivariant output feature $\mathbf{X^{\prime}}$.
+The second module is an SE(3)-equivariant GNN (EGNN) module that, by integrating the processed node feature set $\mathbf{X^{\prime}}$ with the node and edge invariant feature set $\mathbf{Y}$ and $\mathbf{Z}$, outputs the final feature set $\mathbf{F}$.
+The first DGCNN network aggregates features from each point and its neighbors by a local graph, generating richer local geometric representations.
+Before applying the network, we compensate the dominant translational component using ego-velocity and time interval, so that the SO(3)-equivariant layers can focus on rotation-consistent feature learning, and the whole system predicts the residual SE(3) motion on top of the translational initialization.
+A stack of LN [20] network layers is applied in the DGCNN framework.
+LN modifies traditional MLP layers such as Linear, ReLU and Pooling layers for equivariance on semi-simple Lie groups, and preserves SO(3)-equivariance in our case.
+After three iterations, the network outputs the processed node geometric feature set $\mathbf{X^{\prime}}$, whose dimension is lifted to $N\times(C^{\prime}\times 3)$ by LN and retains SO(3)-equivariance with SE(3)-aware pose reasoning.
+The processed node feature set $\mathbf{X^{\prime}}$ of each frame is then passed to the second EGNN module together with physical invariant node feature set $\mathbf{Y}$ and relative invariant edge feature set $\mathbf{Z}$ of this frame for further graph feature aggregation.
+For each node $i$ and its neighbors $\mathcal{N}(i)$ connected by edges, the features are aggregated and updated as follows:
 $\mathbf{m}_{ij}$ $=\phi_{I_{1}}(\exp(\gamma\left\|\mathbf{X}^{{}^{\prime}}_{i}-\mathbf{X}^{{}^{\prime}}_{j}\right\|),\mathbf{Y}_{i},\mathbf{Y}_{j},\mathbf{Z}_{ij}),$ (9)
 $\mathbf{Y}_{i}$ $=\phi_{I_{2}}(\mathbf{Y}_{i},\underset{j\in\mathcal{N}(i)}{\sum}\mathbf{m}_{ij}),$ (10)
-$\mathbf{X}^{{}^{\prime}}_{i}$ $=\mathbf{X}^{{}^{\prime}}_{i}+C\underset{j\in\mathcal{N}(i)}{\sum}(\mathbf{X}^{{}^{\prime}}_{i}-\mathbf{X}^{{}^{\prime}}_{j})\phi_{E}(\mathbf{m}_{ij}),$ (12)
-where $\mathbf{m}_{ij}$ represents the invariant message passed from node $j$ to node $i$ under the Message Passing Neural Network (MPNN) framework [24], $\mathbf{X}^{{}^{\prime}}_{i}$ and $\mathbf{Y}_{i}$ are processed node equivariant and invariant feature of node $i$, respectively.
-$C$ is a constant defined as the number of elements in set $\mathcal{N}(i)$, and $\phi_{I_{1}},\phi_{I_{2}},\phi_{E}$ are MLP networks.
-Differing from the original EGNN that uses relative squared distances $\left\|\mathbf{X}^{{}^{\prime}}_{i}-\mathbf{X}^{{}^{\prime}}_{j}\right\|^{2}$ for invariant feature extraction, inspired by [25], we employ a Gaussian kernel function $\exp(\gamma\left\|\mathbf{X}_{i}^{{}^{\prime}}-\mathbf{X}_{j}^{{}^{\prime}}\right\|)$ with a learnable parameter $\gamma$ to transform the SE(3)-equivariant features $\mathbf{X}$ into SE(3)-invariant representations.
-After four iterations, the network outputs the final node equivariant feature set $\mathbf{X}_{F}=\{\mathbf{X}^{{}^{\prime}}_{i}\}$ and final node invariant feature set $\mathbf{Y}_{F}=\{\mathbf{Y}_{i}\}$ for i=1,2,,Ni=1,2,...,N.
-The final feature $\mathbf{F}$ is defined as the concatenation of $\mathbf{X}_{F}$ and $\mathbf{Y}_{F}$.
+$\mathbf{X}^{{}^{\prime}}_{i}$ $=\mathbf{X}^{{}^{\prime}}_{i}+\frac{1}{\left|\mathcal{N}(i)\right|}\underset{j\in\mathcal{N}(i)}{\sum}(\mathbf{X}^{{}^{\prime}}_{i}-\mathbf{X}^{{}^{\prime}}_{j})\phi_{E}(\mathbf{m}_{ij}),$ (11)
+where $\mathbf{m}_{ij}$ represents the invariant message passed from node $j$ to $i$ under the message passing neural network framework [10], $\mathbf{X}^{{}^{\prime}}_{i}$ and $\mathbf{Y}_{i}$ are processed equivariant and invariant features of node $i$, respectively.
+$\phi_{I_{1}},\phi_{I_{2}},\phi_{E}$ are MLPs.
+Differing from the original EGNN that uses $\left\|\mathbf{X}^{{}^{\prime}}_{i}-\mathbf{X}^{{}^{\prime}}_{j}\right\|^{2}$ for invariant feature extraction, inspired by [36], we employ a Gaussian kernel function $\exp(\gamma\left\|\mathbf{X}_{i}^{{}^{\prime}}-\mathbf{X}_{j}^{{}^{\prime}}\right\|)$ with a learnable parameter $\gamma$ to transform the SO(3)-equivariant features $\mathbf{X^{\prime}}$ into SE(3)-invariant representations.
+After four iterations of the EGNN, the network outputs final node equivariant feature set $\mathbf{X}_{F}=\{\mathbf{X}^{{}^{\prime}}_{i}\}$ and node invariant feature set $\mathbf{Y}_{F}=\{\mathbf{Y}_{i}\}$ for i=1,2,,Ni=1,2,...,N for the processed frame.
+Each frame’s final per-point feature set $\mathbf{F}$ is defined as the concatenation of $\mathbf{X}_{F}$ and $\mathbf{Y}_{F}$.
 III-D2 Transformation Estimation
-In this module, the $\mathbf{L}_{2}$ distances between the final features $\mathbf{F}$ of points in the two point cloud frames are computed.
+In this module, the $\mathbf{L}_{2}$ distances are computed between every feature vector $\mathbf{f}_{s,i}$ and $\mathbf{f}_{t,j}$ in the source and target feature sets $\mathbf{F}_{s}$ and $\mathbf{F}_{t}$, respectively, forming a pairwise-correspondence matrix.
 The $M$ pairs with the smallest feature distances are selected as keypoint correspondences.
-A learnable similarity metric is then defined as:
-$s_{ij}=\exp(-\beta(d_{ij}-\alpha)),$ (13)
-where $s_{ij}$ denotes the similarity between point $i$ in the first frame and point $j$ in the second frame, $d_{ij}$ is their feature distance, and $\alpha,\beta$ are learnable parameters that enhance the robustness of the pipeline to outliers.
-Given the similarity matrix $\mathbf{S}=\{s_{ij}\}_{M\times M}$, we apply the differentiable Sinkhorn normalization algorithm [26] to obtain a soft assignment of correspondences.
-After $K$ iterations, the assigning weight matrix $\mathbf{W}=\{w_{ij}\}_{M\times M}$ is obtained.
-A weighted Singular Value Decomposition (SVD) is then used to estimate the relative transformation $(\mathbf{R},\mathbf{t})$ between the two frames.
+A learnable similarity metric is then defined as: $s_{ij}=\exp(-\beta(d_{ij}-\alpha))$, where $s_{ij}$ denotes the similarity between source point $i$ and target point $j$, $d_{ij}$ is their feature distance, and $\alpha,\beta$ are learnable parameters that enhance the robustness to outliers.
+Given the similarity matrix $\mathbf{S}=\{s_{ij}\}_{M\times M}$, we obtain a soft assigning weight matrix $\mathbf{W}=\{w_{ij}\}_{M\times M}$ of correspondences via the differentiable Sinkhorn algorithm [29].
+A weighted Singular Value Decomposition (SVD) finally estimates the transformation $(\bm{R},\bm{t})$ between the two frames.
 III-D3 Loss Function
-To supervise the transformation estimation, we design a novel loss function that jointly considers rotation, translation, and orientation errors.
-The estimated rotation $\mathbf{R}$ is first converted to Euler angles $\mathbf{E}$, and the loss $\mathcal{L}$ between $(\mathbf{E},\mathbf{t})$ and the ground truth $(\mathbf{E}_{gt},\mathbf{t}_{gt})$ consists of four terms: rotation loss $\mathcal{L}_{r}$, translation loss $\mathcal{L}_{t}$, pitch loss $\mathcal{L}_{p}$, and yaw loss $\mathcal{L}_{y}$:
-$\mathcal{L}_{r}$ $=\left\|\mathbf{E}-\mathbf{E}_{gt}\right\|,$ (14)
-$\mathcal{L}_{t}$ $=\left\|\mathbf{t}-\mathbf{t}_{gt}\right\|,$ (15)
-$\mathcal{L}_{p}$ $=\left\|\mathbf{E}^{pitch}-\mathbf{E}_{gt}^{pitch}\right\|,$ (16)
-$\mathcal{L}_{y}$ $=\left\|\mathbf{E}^{yaw}-\mathbf{E}_{gt}^{yaw}\right\|,$ (17)
+The proposed framework is trained in an end-to-end manner using the ground truth transformation between frames as supervision.
+Specifically, we design a novel loss function that jointly considers rotation, translation, and orientation errors.
+The estimated rotation $\bm{R}$ is first converted to Euler angles $\mathbf{E}$, and the loss $\mathcal{L}$ between $(\mathbf{E},\mathbf{t})$ and the ground truth $(\mathbf{E}_{gt},\mathbf{t}_{gt})$ consists of four terms: rotation loss $\mathcal{L}_{r}$, translation loss $\mathcal{L}_{t}$, pitch loss $\mathcal{L}_{p}$, and yaw loss $\mathcal{L}_{y}$:
+$\mathcal{L}_{r}$ $=\left\|\mathbf{E}-\mathbf{E}_{gt}\right\|,\mathcal{L}_{t}=\left\|\mathbf{t}-\mathbf{t}_{gt}\right\|,$
+$\mathcal{L}_{p}$ $=\left\|\mathbf{E}^{pitch}-\mathbf{E}_{gt}^{pitch}\right\|,\mathcal{L}_{y}=\left\|\mathbf{E}^{yaw}-\mathbf{E}_{gt}^{yaw}\right\|,$ (12)
 where $\mathbf{E}^{pitch}$ and $\mathbf{E}_{gt}^{pitch}$ denote the pitch angles, and $\mathbf{E}^{yaw}$ and $\mathbf{E}_{gt}^{yaw}$ denote the yaw angles, in the estimated and ground truth rotations, respectively.
-The pitch and yaw losses are included to compensate for the relatively low elevation and azimuth resolution of 4D mmWave radar measurements.
+We include additional pitch loss $\mathcal{L}_{p}$ and yaw loss $\mathcal{L}_{y}$ for finer-grained rotation supervision.
+Geometrically, pitch and yaw directly affect forward motion, whereas roll primarily influences lateral and vertical components, which have less impact on forward trajectory error and can be sufficiently constrained by the main rotation loss $\mathcal{L}_{r}$.
 Following [39], we introduce four learnable balancing parameters $s_{r},s_{t},s_{p},s_{y}$ to account for the different units and scales of rotation and translation, leading to the final loss:
@@
 IV-A Datasets
-Two datasets are used for evaluation: the publicly available NTU4DRadLM dataset [6] and a self-collected dataset acquired using a vehicle-mounted 4D mmWave radar and high-precision GPS.
+Two datasets are used for evaluation: the NTU4DRadLM dataset [34] and a self-collected dataset acquired using a vehicle-mounted 4D radar and high-precision GPS.
 IV-A1 NTU4DRadLM Dataset
-The NTU4DRadLM dataset is specifically designed for 4D radar-based localization and mapping.
-It provides accurate ground-truth odometry and includes sequences from both low-speed robot platforms and high-speed unmanned vehicles.
-The dataset was collected using the Oculii Eagle 4D radar, featuring an azimuth resolution of $0.5^{\circ}$, an elevation resolution of $1^{\circ}$, and an operating frequency of 12 Hz.
-NTU4DRadLM contains six splits.
-We use loop1 for training, loop2 for validation, and the remaining four splits for testing.
+This dataset is designed for 4D radar-based localization and mapping.
+It provides ground-truth odometry and includes sequences from both low-speed robot platforms and high-speed vehicles.
+It was collected using the Oculii Eagle 4D radar, which has $0.5^{\circ}$ azimuth and $1^{\circ}$ elevation resolution at 12 Hz operating frequency.
+We use its enhanced radar point cloud and select loop1 for training, loop2 for validation, and the remaining four splits for testing.
 IV-A2 Self-Collected Dataset
@@
 4, the radar was mounted on an SUV equipped with a high-precision GPS system to provide ground-truth odometry.
-The radar has an azimuth resolution of $1.9^{\circ}$, an elevation resolution of $3.5^{\circ}$, and operates at 15 Hz.
-For evaluation, models of all learning-based algorithms trained on the loop1 split of NTU4DRadLM are directly tested on this dataset to compare cross-dataset generalization.
+The radar has $1.9^{\circ}$ azimuth and $3.5^{\circ}$ elevation resolution, operates at 15 Hz.
+All learning-based models trained on NTU4DRadLM are directly tested on this dataset to compare cross-dataset generalization.
 Figure 4: Test vehicle equipped with a 4D radar for the self-collected dataset.
-IV-B Experimental Setup
-IV-B1 Training Details
-The proposed model is trained on a single NVIDIA A100 GPU using the Adam optimizer with a weight decay of $10^{-3}$ for 50 epochs , taking approximately 6 hours.
-The initial learning rate is set to $10^{-4}$ and decays by a factor of 0.1 every 10 epochs.
-The batch size is set at 4.
-IV-B2 Compared algorithms
-For baseline comparison, we select several traditional point cloud registration methods, including Iterative Closest Point (ICP) [27], Normal Distributions Transform (NDT) [28], and Generalized ICP (GICP) [29].
-In addition, we evaluate the State-of-the-Art (SOTA) 4D mmWave radar odometry method Adaptive Probability Distribution-GICP (APDGICP), which serves as the front end of the 4DRadarSLAM framework [2].
-For learning-based methods, we include the self-supervised odometry method RaFlow [13] and the recent supervised method CAO-RONet [15].
-IV-B3 Metrics
-We adopt the evo library [30] to compute the relative translation error ($t_{rel}$) and relative rotation error ($r_{rel}$) for each algorithm, by comparing registered trajectories with the ground-truth.
-Figure 5: Projection of adjacent point clouds onto the corresponding images, where point colors represent compensated node velocities and white circles denote selected keypoints used for matching.
-Figure 6: Comparative odometry results of Equi-RO and baseline algorithms on the NTU4DRadLM dataset for the cp (a), garden (b), nyl (c), and loop3 (d) splits.
-Regions with large rotational motion are highlighted with magnified insets for clarity.
-Figure 7: Translation error (a) and rotation error (b) profiles of Equi-RO and selected baseline algorithms over the full cp split trajectory.
-Figure 8: Comparative odometry results of Equi-RO and baseline algorithms on the self-collected dataset.
-Figure 9: Ablation study results of Equi-RO on the cp (a), garden (b), nyl (c), and loop3 (d) splits of the NTU4DRadLM dataset.
+Figure 5: Projection of adjacent point clouds onto the corresponding images from loop3 split, where point colors represent compensated node velocities and white circles denote selected keypoints used for matching.
+Figure 6: Comparative odometry results of Equi-RO and baselines on the NTU4DRadLM dataset for the cp (a), garden (b), nyl (c), and loop3 (d) splits.
 TABLE I: Quantitative comparison on NTU4DRadLM and Self-collected datasets ($t_{\text{rel}}:\%,r_{\text{rel}}:^{\circ}/\text{m}$)
@@
 APDGICP 4.31 0.0575 3.33 0.0472 3.95 0.0210 4.81 0.0100 4.10 0.0340 10.97 0.0575 34.83 0.0961 22.90 0.0768
+KISS-ICP 4.86 0.0486 4.38 0.0439 4.34 0.0237 3.88 0.0092 4.37 0.0314 11.07 0.0563 25.07 0.0879 18.07 0.0721
+Radar4Motion 4.65 0.0565 3.87 0.0395 4.11 0.0265 4.14 0.0080 4.19 0.0326 13.62 0.0605 22.59 0.0858 18.11 0.0732
 RaFlow 62.12 0.2362 37.07 0.3961 25.93 0.1060 60.31 0.1765 46.86 0.2287 58.43 0.0846 67.59 0.1296 63.01 0.1071
@@
 Equi-RO 3.93 0.0421 3.15 0.0380 4.31 0.0218 3.25 0.0070 3.66 0.0272 8.55 0.0524 10.23 0.0711 9.39 0.0617
-IV-C Experimental Results
-Table I presents a quantitative comparison of our Equi-RO against several baseline algorithms.
-In NTU4DRadLM dataset, our method achieves the lowest translation and rotation errors on the cp, garden, and loop3 splits, and ranks second best on the nyl split.
-Traditional methods such as GICP and APDGICP attain competitive results, while learning-based approaches RaFlow and CAO-RONet exhibit inferior performance.
-This discrepancy is likely due to their design and evaluation primarily on the earlier VoD dataset [31], leading to limited generalization on the newer NTU4DRadLM dataset.
-On average, our algorithm improves translation and rotation accuracy by 10.7% and 20.0%, respectively, compared to the best-performing baseline APDGICP.
-Figure 6 visualizes the estimated trajectories, where our approach (red lines) demonstrates superior accuracy in most splits, particularly in those involving large rotations.
-As presented in Figure 5, we visualize the projection of consecutive point clouds from the loop3 split onto the corresponding images.
-With velocity compensation, the actual motion of dynamic points can be approximated, while the feature extraction algorithm selects keypoints that remain stationary and maintain strong continuity between frames.
-Figure 7 illustrates the translation and rotation errors along the full cp split, excluding the two learning-based algorithms with large errors for clarity.
-While our method does not always yield the lowest translation error over the entire trajectory, it consistently achieves the best rotation accuracy.
-This outcome further validates the effectiveness of our invariant velocity preprocessing, SE(3)-equivariant network design, and adaptive loss function in robustly encoding and estimating rotational motion.
-The comparative results on our self-collected dataset are also summarized in Table I and visualized in Figure 8.
-Due to the lower angular resolution of the employed 4D radar compared to that in the NTU4DRadLM dataset, all methods experience performance degradation.
-Nonetheless, our approach still attains the best translation and rotation accuracy across both splits, despite being trained solely on the NTU4DRadLM dataset, which demonstrates its strong generalization capability.
-IV-D Ablation Study
-To further validate the effectiveness of our proposed algorithm, we conduct an ablation study on the NTU4DRadLM dataset by creating three variant versions of Equi-RO:
-  • Equi-RO with No Invariant Velocity Compensation (NIV): In this variant, node velocities are taken directly from the original Doppler values $v_{i}^{dop}$, and edge velocities are computed as $v_{i}^{dop}-v_{j}^{dop}$, rather than using the compensated counterparts $v_{i}^{dop^{\prime}}$ and $v_{ij}$.
-  • Equi-RO with No Equivariant Network Design (NEN): Here, the LN-based layers are replaced by conventional Linear, ReLU, and Pooling layers, and the EGNN is substituted with a standard MPNN, removing the equivariance property from the network.
-  • Equi-RO with No Adaptive Loss (NAL): In this case, the pitch loss $\mathcal{L}_{p}$ and yaw loss $\mathcal{L}_{y}$ are omitted, and fixed weights ($s_{r}=-8.0$, $s_{t}=-3.0$) are assigned to the rotation and translation losses instead of using learnable balancing parameters.
-The quantitative results of this ablation study are summarized in Table II.
-TABLE II: Ablation study results on the NTU4DRadLM dataset
+Figure 7: Comparative odometry results of Equi-RO and baselines on the two splits of self-collected dataset.
+Figure 8: Comparative odometry results on large-rotation segments of cp (a), garden (b), nyl (c), and loop3 (d) splits of the NTU4DRadLM dataset.
+TABLE II: Quantitative results on large-rotation segments of the NTU4DRadLM dataset
 Method cp garden nyl loop3
 $t_{\text{rel}}$ $r_{\text{rel}}$ $t_{\text{rel}}$ $r_{\text{rel}}$ $t_{\text{rel}}$ $r_{\text{rel}}$ $t_{\text{rel}}$ $r_{\text{rel}}$
-NIV 5.16 0.0542 4.92 0.0601 8.12 0.0432 4.31 0.0088
-NEN 9.60 0.1688 13.27 0.1960 8.97 0.0376 7.97 0.0101
-NAL 5.32 0.0498 4.76 0.0744 9.11 0.0504 5.24 0.0222
+ICP 70.04 0.6056 20.59 0.2884 63.04 0.2001 53.81 0.5660
+NDT Failed Failed 5.76 0.1337 7.23 0.0881 6.84 0.5276
+GICP 16.51 0.5375 4.73 0.1706 8.44 0.0816 6.68 0.5043
+APDGICP 17.61 0.5189 3.63 0.1578 6.83 0.0846 4.94 0.4480
+KISS-ICP 14.35 0.4219 5.33 0.1785 7.23 0.0982 7.23 0.5150
+Radar4Motion 13.26 0.4398 4.86 0.1813 7.38 0.0738 6.52 0.5109
+RaFlow 50.43 2.1401 43.78 0.5364 37.14 0.2808 60.96 0.9989
+CAO-RONet 29.30 1.9730 50.45 0.2681 28.83 0.3657 72.31 1.6608
+Equi-RO 7.23 0.2588 3.38 0.1457 6.47 0.0762 4.64 0.4405
+Figure 9: Ablation study results of Equi-RO on the cp (a), garden (b), nyl (c), and loop3 (d) splits of the NTU4DRadLM dataset.
+Figure 10: Rotation (left) and translation (right) errors under different dynamic-point ratios for Equi-RO and the ablation variant I.
+Points with absolute compensated Doppler velocities greater than 0.1 m/s are treated as dynamic, and numbers on top of bars denote frame counts.
+IV-B Experimental Setup
+IV-B1 Training Details
+We train Equi-RO on a single NVIDIA A100 GPU using Adam with a weight decay of $10^{-3}$ and a batch size of 4 for 50 epochs.
+The initial learning rate is $10^{-4}$ and decays by a factor of 0.1 every 10 epochs.
+For hyperparameters, we set the weight $\lambda=1$ and the number of nearest neighbors $k=6$ in graph construction.
+IV-B2 Compared Algorithms
+We first select several traditional registration methods ICP, NDT, and GICP as baselines.
+We also evaluate an ICP-based method KISS-ICP [31], a 4D radar odometry method Radar4Motion [16], and Adaptive Probability Distribution-GICP (APDGICP) [35], the State-of-the-Art (SOTA) 4D radar odometry method on NTU4DRadLM dataset.
+For learning-based methods, we include RaFlow [7] and CAO-RONet [19].
+These two methods are also trained on the loop1 split of NTU4DRadLM using their default settings.
+IV-B3 Metrics
+We adopt the evo library [11] to compute the relative translation error ($t_{rel}$) and relative rotation error ($r_{rel}$) for each algorithm, by comparing registered trajectories with the ground truth.
+We average errors over 100 equal segments (with $\delta=L/100$) of each trajectory.
+IV-C Experimental Results
+Fig.
+5 qualitatively shows that velocity compensation can help estimate the actual motion of dynamic points, while the feature extraction can select keypoints that remain stationary and consistent between frames.
+We present quantitative comparison of our Equi-RO against baseline algorithms in Table I.
+In the NTU4DRadLM dataset, our method achieves the lowest translation and rotation errors on the cp, garden, and loop3 splits, and ranks second best rotation error on the nyl split.
+GICP, APDGICP, KISS-ICP and Radar4Motion attain competitive results, while learning-based approaches RaFlow and CAO-RONet exhibit inferior performance.
+This discrepancy is likely because they are originally developed for the earlier VoD dataset [24], and appear less suitable to the newer NTU4DRadLM dataset.
+On average, our algorithm improves translation and rotation accuracy by 10.7% and 13.4%, respectively, compared to the best baselines.
+Fig.
+6 visualizes the estimated trajectories, where our method (red lines) demonstrates superior accuracy in most splits.
+The comparative results on self-collected dataset are also shown in Table I and Fig.
+7.
+All methods degrade on this dataset with lower radar resolution.
+Nonetheless, our method still attains the best translation and rotation accuracy across both splits, despite being trained solely on the NTU4DRadLM dataset, demonstrating its strong generalization capability.
+We further evaluate performance on sharp-turn segments from the NTU4DRadLM dataset.
+Equi-RO consistently achieves the best or second-best accuracy, as shown in Fig.
+8 and Table II.
+These results confirm that the proposed equivariant and invariant features, and the equivariant network design are especially effective in scenarios with large rotations.
+IV-D Ablation Study
+We conduct an ablation study on the NTU4DRadLM dataset by creating five variants of Equi-RO:
+(I) w/o invariant velocity compensation: In this variant, node velocities are set to the original Doppler values $v_{i}^{dop}$, and edge velocities are computed as $v_{i}^{dop}-v_{j}^{dop}$, rather than using the compensated counterparts $v_{i}^{dop^{\prime}}$ and $v_{ij}$.
+(II) w/o equivariant network design: Here, the LN-based layers are replaced by conventional Linear, ReLU, and Pooling layers, and the EGNN is substituted with a standard message passing neural network, removing the equivariance property.
+(III) w/o adaptive loss: In this case, the pitch loss $\mathcal{L}_{p}$ and yaw loss $\mathcal{L}_{y}$ are omitted, and fixed weights ($s_{r}=-8.0$, $s_{t}=-3.0$) are assigned to the rotation and translation losses instead of learnable parameters.
+(IV) w/o Doppler velocity: The compensated node velocity $v_{i}^{dop^{\prime}}$ and edge velocity $v_{ij}$ are set to zero in this case.
+(V) w/o initial translation: In this ablation, we remove the initial translation based on ego-velocity.
+The quantitative results are summarized in Table III.
+TABLE III: Ablation study results on the NTU4DRadLM dataset
+Method cp garden nyl loop3
+$t_{\text{rel}}$ $r_{\text{rel}}$ $t_{\text{rel}}$ $r_{\text{rel}}$ $t_{\text{rel}}$ $r_{\text{rel}}$ $t_{\text{rel}}$ $r_{\text{rel}}$
+I 5.16 0.0542 4.92 0.0601 8.12 0.0432 4.31 0.0088
+II 9.60 0.1688 13.27 0.1960 8.97 0.0376 7.97 0.0101
+III 5.32 0.0498 4.76 0.0744 9.11 0.0504 5.24 0.0222
+IV 5.25 0.0523 4.94 0.0614 8.67 0.0505 6.78 0.0109
+V 5.21 0.0514 3.97 0.0427 6.87 0.0401 4.05 0.0095
 Equi-RO 3.93 0.0421 3.15 0.0380 4.31 0.0218 3.25 0.0070
@@
 These trajectories demonstrate that the full Equi-RO model consistently achieves the best performance.
-All three ablated variants show notable performance degradation, particularly in scenarios involving large rotations.
-The NEN variant, which discards equivariant network design, exhibits the worst results across most splits, underscoring the critical importance of equivariant architecture in our framework.
+All five ablated variants show notable performance degradation.
+The variant II, which discards equivariant network design, exhibits the worst results across most splits, underscoring the critical importance of equivariant architecture in our framework.
+As shown in Fig.
+10, we compare the per-frame average translation and rotation errors of the Equi-RO method with the no invariant velocity compensation variant I across different dynamic-point ratios.
+As the ratio of dynamic points increases, the errors of the variant I consistently grow, while the full method remains relatively stable.
+This demonstrates that compensated invariant Doppler features effectively improve robustness in dynamic scenes.
 IV-E Limitations
-Although Equi-RO demonstrates superior accuracy and robustness, its performance is affected by the relatively low resolution of radars, as shown in the self-collected dataset.
-Additionally, while our approach meets general real-time requirements with a processing speed of 76.75 ms per frame on a single NVIDIA A100 GPU, further algorithmic optimization is necessary for deployment on on-board platforms with stricter latency budgets.
+Although Equi-RO demonstrates superior accuracy and robustness, its performance is still affected by the relatively low resolution of radars, as shown in the self-collected dataset.
+Moreover, while our method achieves real-time performance at 58.15 ms per frame on a single NVIDIA A100 GPU, further optimization is necessary for deployment on on-board platforms.
+Efficiency can be improved via C++ migration and TensorRT acceleration, bridging the gap between the research prototype and mobile platform constraints.
 V Conclusion
@@
 This work highlights the potential of equivariant network design for radar-based perception and provides a new perspective for point cloud processing in robotics and autonomous systems.
-Future work will extend the framework to multi-modal fusion with IMUs and cameras and explore its application to other 3D point clouds with physical attributes.
+Future work will extend the framework to multi-modal fusion with IMUs and cameras, and explore its application to other 3D point clouds with physical attributes.